OpenStudy (anonymous):

Given A (-6,4) and M(-1, -1) find B. 1) (-4, 6) 2) (-6, -4) 3) (6,4) 4) (4, -6)

OpenStudy (anonymous):

is B the vector from A to M ?

OpenStudy (anonymous):

B (x,y) (-6+x)/2=-1, x=4 (4+y)/2=-1, y=-6